OpenStudy (leahmarieed):

Solve the following system of equations: x − 2y = 6 2x − 4y = 10 Infinitely many solutions No solutions (0, 0) (6, 10)

OpenStudy (welshfella):

first step: multiply the first equation by 2 what do you get ? just multiply each term by 2

OpenStudy (leahmarieed):

X-4Y=12?

OpenStudy (welshfella):

you haven't multiplied the x by 2

OpenStudy (leahmarieed):

x2

OpenStudy (welshfella):

2x giving 2x - 4y = 12 so we have 2 equations 2x - 4y = 12 and 2x - 4y = 10 - do you notice anything odd about these 2 equations?

OpenStudy (welshfella):

the 2 left sides are the same but the right ones are different

OpenStudy (leahmarieed):

ok

OpenStudy (welshfella):

so are there any values of x and y which will fit the 2 equations?

OpenStudy (leahmarieed):

2

OpenStudy (welshfella):

what 2 values?

OpenStudy (leahmarieed):

6,10

OpenStudy (welshfella):

x = 6 and y = 10?

OpenStudy (leahmarieed):

yes

OpenStudy (welshfella):

no try first one 2x - 4y = 12 2(6) - 4(10) = 12 - 40 = -28 NOT 12

OpenStudy (leahmarieed):

ohhhh

OpenStudy (leahmarieed):

so would there be no solutions

OpenStudy (welshfella):

look at the 2 equations again can the left side be equal to 10 ANd 12?

OpenStudy (welshfella):

thats correct its impossible for LHS to be equal to 10 and 12 so NO SOLUTIONS

OpenStudy (welshfella):

if we subtract the 2 equations we get 0 = 10 0 =12 which is of course nonsense

OpenStudy (welshfella):

Note that by multiplying equation by a constant number does not change the equation. Its basically the same equation .

OpenStudy (leahmarieed):

i see, thank you so much

OpenStudy (welshfella):

if we drew the graph of x - 2y = 6 and 2x - 4y = 12 its would be the same straight line.

OpenStudy (welshfella):

yw